OpenStudy (anonymous):

The value of y varies directly with x, and y = 1 when x = 2. Find y when x = 4.

OpenStudy (solomonzelman):

The value of y varies directly with x y=kx

OpenStudy (solomonzelman):

y = 1 when x = 2 1=k2

OpenStudy (solomonzelman):

1=k 2 (times, not square) k= 1/2

OpenStudy (solomonzelman):

Find y when x = 4 y=(1/2)x y=(1/2) 4 y=2